(no subject)

John Shields (jshields@ksera)
Wed, 7 Jul 1999 00:02:31 -0500

Shift Summary for eve shift, 07/06/99

On Shift: J. Shields, E. Turner

Hours of beam available at all intensities: 6.4
Hours of beam available at nominal intensity: 2.5
Hours of E832 data (special runs not included): 6.4
Hours of special runs (laser scans, muon runs...): 0.2 (cosmic)

Accounting of lost time:

tevatron MI problems result in sporatic and reduced-intensity beam, access
to replace HA power supply, replacement of faulty tape drive

========================================================================
Runs ended this shift:
Run 13827 - BEAM_832
Run 13829 - cosmic

Tapes brought to FCC this shift:
(including tapes for online split)

QKS092-094, QKC092-094, QKK092-094, QKB032

online split: QKLA06

========================================================================
Accesses during shift: 1 (but a very prolonged access, unfortunately)

21:10-23:00 SET and EB valiantly attempt to repair a HA power supply

========================================================================
Overview of shift:

E832 datataking with high and low quality beam, tape drive is repaired,
SET and EB go on access

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13827 16:25 swapped blue ice on alcohol bubbler. temp ~4.5

18:05 MCR calls to tell us to expect variable beam, due
to tuning problems w. MI

20:05 SDAQ complains that s9-UCLA-PS4 has gone bad, this
turns out to correspond to HA, etc

21:10 SET and EB go on access and discover a blown fuse on
+5V power supply. However, +5V happily supplies power
while -5V does not.

21:30 Install new geometry constants. Changed DC-TDC intime
window by -192 counts.

21:50 tapes to Feynman Center

21:57 Rick Coleman repairs the tape drives on ktev2.
ktev2 is shut down briefly.

13831 22:35 swapped blue ice on alcohol bubbler. temp ~5.0
conclusion: this may now be unnecessary.

23:00 SET,EB out of access w. power supply, decide to run w/o
HA while it undergoes repairs. Currently suspect that
fuses were labeled incorrectly. Hence, the blown fuse
does not appear to correspond to the bad -5V supply.

23:10 beam is restored

24:00 SET, EB go back on access to replace the now-repaired
power supply

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

1) don't panic about low CO2. Dave N. is aware of the situation and will
swap it first thing in the morning

2) keep an eye on the alcohol temperature. The blue ice may or may not
still be necessary, now that the weather is a bit cooler.